OpenStudy (christos):

Please help me solve 4^x - 2^x - 12 = 0

jimthompson5910 (jim_thompson5910):

Hint: Let z = 2^x so z^2 = (2^x)^2 ---> z^2 = 4^x

jimthompson5910 (jim_thompson5910):

which means you can turn 4^x - 2^x - 12 = 0 into z^2 - z - 12 = 0

OpenStudy (anonymous):

x = 2

OpenStudy (christos):

x = 6 or x = 2 ? I got x = 6 is my result correct?

jimthompson5910 (jim_thompson5910):

you would then solve for z once you get your two solutions for z, you would use them to find the solution for x

jimthompson5910 (jim_thompson5910):

when you solve z^2 - z - 12 = 0 for z, you get z = 4 z = -3

jimthompson5910 (jim_thompson5910):

ignore the -3 since 2^x is always positive

jimthompson5910 (jim_thompson5910):

so z = 2^x 4 = 2^x x = ???

OpenStudy (christos):

2 :(

jimthompson5910 (jim_thompson5910):

yep